A bicyclist rode into the country for 5 hours. In returning, her speed was 5 mi/h faster and the trip took 4 hours. What was her speed each way?
Answer by Paul(988)

His going speed = x
His returning speed = x+5

Equation:
4(x+5)=5(x)
4x+20=5x
x=20

20+5=25

Hence, his going speed was 20mph and his returning speed was 25mph.
Paul.